### v1

# USE-LFA: A Data-Driven Framework for Urban Air Mobility Site Evaluation using Latent Factor Analysis

Abstract
This study introduces the Urban Site Evaluation using Latent Factor Analysis (USE-LFA) framework, a data-driven methodology for selecting optimal urban port locations, demonstrated through a case study in Seoul. USE-LFA identifies six critical urban factors categorized into suitability (operational feasibility) and attractiveness (operational merits), enabling a holistic evaluation of potential sites. The framework employs latent factor analysis to uncover hidden patterns in urban data, integrating diverse attributes such as floating population, economic activity, and infrastructure connectivity. A composite metric, the v-score, balances suitability and attractiveness, offering flexibility to prioritize specific strategic goals. Spatial analysis reveals distinct regional typologies—balanced, suitability-biased, and attractiveness-biased regions—highlighting the adaptability of USE-LFA across varying urban contexts. Beyond its application to vertiports, this versatile framework provides valuable insights for integrating emerging mobility technologies and transit hubs into complex urban environments, supporting sustainable urban planning and informed policymaking.
1 Introduction
The advent of Urban Air Mobility (UAM) signifies a paradigm shift in urban transportation, offering promising applications such as air taxis, air ambulances, and air shuttles (EASA, 2021; Goyal et al., 2018). UAM holds great potential to reduce congestion, improve emergency response times, and alleviate environmental pollution (Yedavalli and Mooberry, 2019). Ensuring its sustainability requires not only environmentally efficient aircraft designs that minimize emissions and noise pollution (Afonso et al., 2021) but also the strategic placement of vertiports--facilities where UAM aircraft can takeoff, land, and be serviced. The vertiport site selection problem addresses this multifaceted challenge that requires comprehensive consideration of various urban factors to ensure effective, efficient, and sustainable integration of UAM systems into urban environments (Holden and Goel, 2016; Cohen et al., 2021; Garrow et al., 2021; Yoon et al., 2025).
In the nascent stages of UAM system development, vertiport site selection is particularly important. Urban areas are expected to undergo significant changes with the introduction of UAM and vertiports, necessitating a comprehensive approach that takes into account for various urban factors such as operational, social and economic factors. The Community Integration Working Group (CIWG) highlights extensive considerations on the integration of UAM into urban environments, and NASA has organized these insights into a comprehensive list of 18 main themes of considerations for vertiport placement (Mendonca et al., 2022). These themes underscore the multifaceted nature of the challenge, highlighting the need for a comprehensive framework that addresses various considerations.
Existing studies on vertiport site selection have employed various methodologies such as clustering algorithms, simulation-based methods, custom optimization models, expert surveys, and GIS-based approaches. These approaches have provided valuable insights and laid a strong foundation for understanding this complex problem. Nevertheless, given the intricate dynamics of urban environments, there remains an opportunity to expand upon these efforts by incorporating more comprehensive perspectives on urban structures and dynamics.
To contribute to this growing body of work, we propose a novel framework leveraging Latent Factor Analysis (LFA) on regional data to uncover latent factors that characterize the urban environment for vertiport integration. Our approach introduces the concept of location-signatures, the structural latent factors that constitute urban environments, grouped into two pillars: suitability and attractiveness.
  • Suitability refers to the operational feasibility of a location for a vertiport. Factors such as community acceptance and the potential for minority gentrification are critical considerations for suitability. A region may be deemed unsuitable if there is strong community opposition or if the region is overly affluence, posing social justice barriers.
  • Attractiveness pertains to the operational merits of the vertiport. A region is deemed attractive if it is expected to capture significant existing demand or has operational advantages. Factors such as potential connectivity to nearby public transit systems or high economic activity contribute to a location’s attractiveness for a vertiport.
In essence, suitability is a “must-have” - a basic requirement for vertiport placement, while attractiveness is a “nice-to-have” - factors that enhance the potential success and integration of the vertiport. Together, these pillars form the foundation for effective vertiport site selection in complex urban settings.
LFA emerges as a critical solution, uncovering and quantifying latent location-signatures that reflect underlying urban structures and dynamics. This approach enables a comprehensive evaluation that integrates social considerations like community dynamics with economic factors such as economic activeness. By consolidating these various factors into a composite v-score metric, LFA offers a holistic perspective on urban environments, ensuring sustainable vertiport integration that balances operational feasibility with operational merits.
This paper is structured as follows: Section 2 provides details on existing work with diverse methodologies. Section 3 introduces the framework and methodology. Section 4 presents data and results for Seoul case study, and Section 5 discussion furhter result analysis. Finally, Section 6 concludes the paper with suggestions for future studies.
2 Literature Review
Numerous studies have explored the problem of vertiport site selection using various methodologies. Expanding on the comprehensive literature review by Long et al. (2023), we further identified various approaches in three groups: demand estimation, expert opinions, and geographic information system approach.
2.1 Demand Estimation
Clustering algorithms, particularly K-means clustering, are widely used to identify vertiport locations. Tarafdar et al. (2021) expanded on the work by Rimjha et al. (2020) and used the K-means algorithm and center-of-mass approaches for UAM landing site selection. Rajendran and Zack (2019) also applied K-means clustering on the predicted mobility data from the NYC Taxi and Limousine Commission data for the strategic placement of the vertiports in New York City. Bulusu et al. (2021) developed a traffic demand analysis method to identify areas with a high potential for UAM demand and used iterative K-means clustering to select the vertiport sites in the San Francisco Bay Area.
Simulation-based methods involve tools such as MATSim, aiming to provide insights into UAM demand prediction and identify areas with high potential demand as vertiport locations. Articles by Rothfeld et al. (2018) and Balac et al. (2019) introduced an approach utilizing agent-based simulations to assess the demand for UAM and determine optimal vertiport sites with travel patterns, user behaviors, and infrastructure constraints. Wu and Zhang (2021) used the simulation output of the Tampa Bay Regional Planning Model (TBRPM) along with socio-demographic data to explore interactions between UAM travel demand and vertiport locations. Building upon this work, Zhao et al. (2022) examined the environmental impacts of UAM implementation in the Tampa Bay area to provide insights into sustainable vertiport planning and location optimization.
Optimization models are employed by many researchers, tailored to predict UAM demand and optimize vertiport locations. Ploetner et al. (2020) predicted UAM demand in Europe based on possible adoption rates of UAM, incorporating savings in travel time, willingness to pay, and operational constraints to optimize vertiport locations that maximize demand coverage. Rath and Chow (2022) focused on NYC Taxi users, developing a p-hub median location problem that maximizes travel time savings to multiple airports while incorporating factors such as user choice behavior and demand elasticity. Chae et al. (2023) identified vertiport locations based on geographic and urban constraints, analyzing population density, commuter patterns, and accessibility to maximize demand coverage and ensure equitable access across the region.
2.2 Expert surveys and opinions
The expert surveys and opinions are employed through methods like the Analytical Hierarchy Process (AHP) and qualitative evaluation to weight various criteria for vertiport site selection. Fadhil (2018) utilized a two-step AHP-Delphi method, engaging 15 experts to evaluate and weight the socioeconomic factors influencing site selection. Similarly, Lee et al. (2023) utilized the AHP method in their studies. They integrated aviation expert knowledge into spatial data analysis, addressing airspace regulations, obstacle clearance requirements, and noise considerations for UAM route design and vertiport placement. Preis (2021) incorporated expert interviews and workshop discussions to formulate criteria for vertiport site selection and anticipate effects. Nordstrom (2022) combined expert knowledge with public surveys and analytical approaches to identify optimal vertiport locations. Furthermore, Cho et al. (2022) explored UAM’s potential in Seoul by analyzing socio-demographic factors and public adoption readiness. Their study integrated expert insights to address challenges such as safety, noise, and infrastructure requirements.
2.3 Geographic information system (GIS) approach
Geographical information system (GIS) approach is utilized to integrate multiple urban factors into spatial analysis for vertiport site selection. Wu and Zhang (2021) incorporated physical and regulatory factors for UAM operations through a three-dimensional GIS map from lidar data to identify candidate locations in Florida. Sinha and Rajendran (2022) expanded on Rajendran and Zack (2019), employing a multi-criteria warm start technique that considers socioeconomic variables as layers and utilized the multi-criteria decision-making technique to compute the total score of suggested sites. Sheth (2023) developed a regional modeling tool named VAMOS!, which utilizes numerical score inputs from users for vertiport placement, including various attributes like zoning, environmental impact, and intermodal transportation systems as layers. Wei et al. (2023) conducted a case study on vertiport placement in the San Francisco Bay Area, developing a systematic framework using GIS. Their study identified three key parameters--safety, access, and equity--to evaluate site suitability. Additionally, Yoon et al. (2025) proposed a strategic GIS-based methodology to integrate UAM with highway infrastructure in the Seoul Metropolitan Area, evaluating 148 candidates based on geographic information, origin-destination volume, and travel time to identify 56 optimal vertiport locations.
Many studies primarily prioritized demand estimation and economic aspects, neglecting the complex nature of UAM integration in urban environments. Expert-driven analyses, while providing valuable insight, are particularly restrictive given the nascent stage of vertiport development and the evolving nature of UAM systems. Although GIS approaches are capable of accounting for different dimensions within the urban area, these studies often analyze data independently, overlooking the interconnected interplay of urban factors.
3 Methodology
We introduce USE-LFA, a novel framework for UAM site evaluation framework using latent factor analysis (LFA) that encapsulates multiple urban considerations into suitability and attractiveness. USE-LFA is employed on regional data to identify latent factors that characterize urban environments relevant to vertiport integration (see Figure 1). This study employs principal axis factoring (PAF) with Varimax rotation and regression-based scoring to identify latent factors underlying urban attributes across regions.
Figure 1: Flow of VESLA, from the city-specific urban attributes to the suitability and attractiveness scores for the regions.
3.1 Input Data Preparation for VESLA
The table below provides an example of data attributes that can be used as input for VESLA. The vertiport-related themes are predetermined and relevant data is gathered accordingly. The input attributes are flexible and can be tailored to suit the unique characteristics of each study area. The description column outlines the normalization process, an essential step for LFA to address variations in census tract sizes and population densities across the regions.
Table 1: Input Description of the Urban Attributes
Theme N Urban Attributes Data Description
Population 1 Housing Density Count of housing units per km2 (units/km2).
2 Floating Population Hourly floating population (8 AM–8 PM) density (pax/km2).
3 Foreign Population Hourly foreign floating population density (pax/km2).
4 Commuting Population Ratio of morning commute flows to the registered population.
5 Incoming Population Ratio of inbound to outbound commuters during peak hours.
6 Disabled Population Percentage of disabled residents relative to the total population (%).
7 Beneficiary Welfare recipients as a percentage of the census tract population (%).
8 Senior Population Proportion of elderly residents (65+) in the registered population (%).
Socioeconomic 9 Average Income Monthly average income aggregated at the census level (KRW/month).
10 Spending Ratio Total expenditures as a percentage of monthly average income (%).
11 Transportation Expenditure Transit costs (bus, express bus, taxi, train, and metro) as a percentage of monthly income (%).
Transport Infrastructure 12 Bus Stop Density Total bus stops per km2, calculated via spatial joins with administrative boundaries (stops/km2).
13 Metro Coverage Percentage of land within a 1,000m buffer around metro stations (%).
14 Metro Users Per Station Daily metro users per station in a census tract during weekdays (users/station).
15 Bus Users Per Stop Daily bus boardings per bus stop in a census tract during weekdays (users/stop).
Land Use and Facilities 16 Parking Space Density Parking spaces per km2 based on POI data (spaces/km2).
17 Touristic Facility Density Count of touristic facilities per km2 from POI data (facilities/km2).
18 Cultural Facility Density Cultural facilities per km2, spatially aggregated (facilities/km2).
19 Energy Usage Intensity Annual energy consumption normalized by area (kWh/km2/year).
20 Employment Density Total employees per km2 based on business registries (employees/km2).
21 Land Price Land value per km2 from official surveys (KRW/km2).
Spatial Proximity 22 Distance to Helipads Euclidean distance from census centroid to the nearest helipad location (km).
23 Distance to Hospitals Euclidean distance from census centroid to the nearest hospital location (km).
24 Distance to VFR Routes Euclidean distance from census centroid to the nearest VFR route (km).
25 Distance to Fire Stations Euclidean distance from census centroid to the nearest fire station (km).
3.2 Z-Score Standardization
Due to the framework’s sensitivity to data scale, Z-score standardization was applied to the input data. This process ensures that attributes with larger variances do not overshadow others, enabling a more balanced and accurate analysis.
3.3 Latent Factor Analysis (LFA)
The core constitution of VESLA is a robust mathematical model that decomposes complex regional data into interpretable factors. Let $\mathbf{A}\in\mathbb{R}^{N\times R}$ represent the standardized urban attribute matrix where $N$ is the number of standardized urban attributes (indexed by $i=1,2,\dots,N$) and $R$ is the number of regions in the study area (indexed by $j=1,2,\dots,R$). Each element $a_{ij}$ encodes the standardized measurement of urban attribute $i$ in region $j$.
The foundational relationship between the urban attributes and their underlying latent factors is formalized as:
$\boldsymbol{A}=\mathbf{L}\mathbf{F}+\mathbf{E}.$ (1)
Where $\mathbf{L}\in\mathbb{R}^{N\times M}$ represents the urban attribute loading matrix with elements $\ell_{im}$ capturing the intensity of relationships between urban attribute $i$ and latent factor $m$. The factor score matrix $\mathbf{F}\in\mathbb{R}^{M\times R}$ contains elements $f_{mj}$ indicating the degree to which region $j$ embodies latent urban factor $m$. The matrix product $\mathbf{LF}$ can be decomposed as $\sum_{m=1}^{M}\boldsymbol{\ell}_{m}\mathbf{f}_{m}$, where $\boldsymbol{\ell}_{m}=[\ell_{1m},\ldots,\ell_{Nm}]^{\top}\in\mathbb{R}^{N}$ represents the loading vector for factor $m$ across all attributes, and $\mathbf{f}_{m}=[f_{m1},\ldots,f_{mR}]\in\mathbb{R}^{R}$ represents the score vector for factor $m$ across all regions. Finally, $\mathbf{E}\in\mathbb{R}^{N\times R}$ captures region-specific variance unique to each attribute. This matrix factorization approach enables dimensionality reduction while preserving the essential structure of urban environments, allowing meaningful latent factors to be discovered from regional input urban attributes.
Initial Communality Estimation
Principal Axis Factoring (PAF) is used to extract the urban latent factors by isolating the shared variance among the urban attributes. The process undergoes an iterative procedure that refines estimates of communalities--the portion of each attribute’s variance explained by the latent factors.
For standardized input urban attribute matrix $\mathbf{A}$, the covariance structure reduces to:
$\text{Cov}(\boldsymbol{A}_{standardized})=\text{Corr}(\boldsymbol{A}_{raw})=\mathbf{R},$ (2)
where $\mathbf{R}$ hereby indicates the correlation matrix of the original (non-standardized) urban attributes. This equivalence enables the direct interpretation of factor loadings as standardized coefficients and scale-invariant comparison of urban attribute relationships. On correlation matrix $\mathbf{R}$, the initial communality $\hat{\mathbf{c}}_{i}^{2(0)}$ for urban attribute $i$ is estimated via squared multiple correlations:
$\hat{\mathbf{c}}_{i}^{2(0)}=1-{\frac{1}{r^{ii}}},$ (3)
where $r^{ii}$ is the $i^{th}$ diagonal element of the inverse correlation matrix $\mathbf{R^{-1}}$. This initialization step approximates how much variance in urban attribute $i$ is explained by its linear relationship with other attributes.
Iterative Procedure
Adjusted Correlation Matrix
The iterative factor extraction process subsequently operates on an adjusted correlation matrix $\mathbf{R}^{*}$, where diagonal elements (original correlations of 1) are replaced with communality estimates to focus on shared variance among the urban attributes:
$\forall i:\mathbf{R^{*}}=\mathbf{R}-\text{diag}(\mathbf{I}_{i}-\mathbf{\hat{c}}_{i}^{2(0)}).$ (4)
Eigenvalue Decomposition
The adjusted correlation matrix $\mathbf{R}^{*}$ is decomposed via eigenvalue decomposition:
$\mathbf{R}^{*}=\mathbf{Q}\mathbf{\Lambda}\mathbf{Q}^{\top}\quad\text{with}\quad\begin{cases}\mathbf{Q}=[\mathbf{q}_{1}|\cdots|\mathbf{q}_{N}]\in\mathbb{R}^{N\times N}\\ \mathbf{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{N})\end{cases}$ (5)
where $\mathbf{Q}$ represents the eigenvectors with $\mathbf{q}_{m}=[q_{1m},...,q_{Nm}]^{\top}\in\mathbb{R}^{N}$ and $\mathbf{\Lambda}$ contains the corresponding eigenvalues.
Factor Loading Estimation
At each iteration $k$, the urban factor loadings are computed as follows:
$\mathbf{L}^{(k)}=\mathbf{Q}_{N}^{(k)}\left(\mathbf{\Lambda}_{N}^{(k)}\right)^{1/2}\quad\in\mathbb{R}^{N\times N}$ (6)
where $\mathbf{Q}_{N}^{(k)}$ contains the eigenvectors and $\mathbf{\Lambda}_{N}^{(k)}$ the corresponding eigenvalues at iteration $k$.
Updating Communalities
Using the current loading matrix $\mathbf{L}^{(k)}$, updated communalities are calculated as:
$\hat{c}_{i}^{2(k+1)}=\sum_{m=1}^{M}\left(\ell_{im}^{(k)}\right)^{2}$ (7)
where $\ell_{im}^{(k)}$ denotes the loading of the urban attribute $i$ on factor $m$ at iteration $k$. These values represent the cumulative explained variance for each attribute across all latent urban factors. The updated communalities $\hat{c}_{i}^{2(k+1)}$ then replace the diagonal entries of the correlation matrix, generating the adjusted matrix $\mathbf{R}^{*(k+1)}$ for the subsequent iteration $k+1$.
Convergence Criterion
The iteration terminates when the total communality change reaches the threshold:
$\sum_{i=1}^{N}\left|\hat{c}_{i}^{2(k+1)}-\hat{c}_{i}^{2(k)}\right|<\epsilon\quad(\epsilon=10^{-5})$ (8)
Final factor loadings $\mathbf{L}^{*}$ are obtained at the convergence iteration $k_{\text{final}}$. Following the Kaiser criterion, only factors with eigenvalues $\lambda_{m}\geq 1$ are retained, ensuring each factor explains at least as much variance as a single original attribute.
The urban factor loadings in $\mathbf{L}^{*}$ can be interpreted as correlation coefficients between each urban attribute and the extracted factors, with larger absolute values indicating stronger relationships. These loadings are essential for identifying which urban characteristics cluster together to form meaningful latent dimensions of urban environment.
Factor Rotation
To enhance interpretability, the Varimax rotation is implemented on the factor structure. The Varimax rotation maximizes the sum of variances of squared loadings within each factor, which tends to produce a simpler, more interpretable factor structure with loadings closer to either 0 or 1. The rotated factor loading matrix is given by:
$\mathbf{L}^{R}=\mathbf{L}^{*}\mathbf{V},$ (9)
where $\mathbf{V}\in\mathbb{R}^{M\times M}$ denotes the orthogonal Varimax rotation matrix satisfying $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{M}$.
Factor Score Estimation
Regional urban factor scores were computed via the regression estimator:
$\mathbf{B}=\Bigl{(}\mathbf{L}^{\top}\mathbf{R}^{-1}\mathbf{L}\Bigr{)}^{-1}\mathbf{L}^{\top}\mathbf{R}^{-1}\quad\in\mathbb{R}^{M\times N},$ (10)
where $\mathbf{R}\in\mathbb{R}^{N\times N}$ is the original correlation matrix and $\mathbf{L}$ refers to the rotated factor loading matrix $\mathbf{L}^{R}$. The factor score matrix $\mathbf{F}\in\mathbb{R}^{M\times R}$ is then obtained through:
$\mathbf{F}=\mathbf{B}\mathbf{A}=\left[\sum_{i=1}^{N}\beta_{mi}a_{ij}\right],$ (11)
with component-wise form:
$f_{mj}=\sum_{i=1}^{N}\beta_{mi}a_{ij}\quad\forall m\in{1,\ldots,M},\ j\in{1,\ldots,R}.$ (12)
$a_{ij}\in\mathbb{R}$ represents the standardized value of urban attribute $i$ in region $j$. The summation aggregates the contributions from all $N$ urban attributes in region $j$, each weighted by the corresponding $\beta$ coefficient, to yield latent factor scores for a region.
The factor score is then divided into suitability and attractiveness, through the following composite definition matrix $\mathbf{T}\in\{0,1\}$:
$\mathbf{T}=\begin{pmatrix}t_{11}&\cdots&t_{1M}\\ t_{21}&\cdots&t_{2M}\end{pmatrix}$ (13)
where $t_{1m}=1$ if factor $m\in\text{Suitability}$, $t_{2m}=1$ if factor $m\in\text{Attractiveness}$, with all other entries being 00.
The composite score matrix $\hat{\mathbf{F}}$ is computed through:
$\hat{\mathbf{F}}=\mathbf{T}\mathbf{F}=\begin{pmatrix}\sum_{m\in\mathcal{S}}f_{m1}&\cdots&\sum_{m\in\mathcal{S}}f_{mR}\\ \sum_{m\in\mathcal{A}}f_{m1}&\cdots&\sum_{m\in\mathcal{A}}f_{mR}\end{pmatrix}\quad\in\mathbb{R}^{2\times R},$ (14)
where $\mathcal{S}$ = Suitability factor indices and $\mathcal{A}$ = Attractiveness factor indices, such that $\mathcal{S,A}\subset\{1,\dots,m\}$ and $\mathcal{S}\cap\mathcal{A}=\emptyset$.
The framework distills complex urban attributes into interpretable dimensions by employing factor analysis to identify latent urban factors, and then strategically categorizing these factors into distinct Suitability and Attractiveness. This binary categorization offers a structured analytics of a dual-perspective evaluation of urban regions while acknowledging the multidimensional nature of urban environment. The framework’s mathematical formalization ensures reproducibility across different urban contexts, while its flexibility allows researchers and planners to adapt factor classifications based on specific policy objectives or theoretical considerations.
4 Case study in Seoul
The framework was demonstrated for Seoul as a case study. Globally recognized as one of the most UAM-ready regions (KPMG, 2023), Seoul has an extensive helipad infrastructure and has been the subject of numerous studies examining its potential for UAM implementation (Cho et al., 2022; Hwang and Hong, 2023; Yoon et al., 2025).
4.1 Data Description
Our analysis incorporated diverse data from Seoul to ensure a comprehensive evaluation. NASA has reportedly suggested 18 groups of themes with more than 450 considerations for vertiport placement (Mendonca et al., 2022). From these, 25 relevant regional urban attributes were selected for the vertiport site selection in Seoul (see Table 2). The selected data was obtained from the Seoul Big Data Campus and the Seoul Open Data Plaza (Seoul Metropolitan Government, 2024a, b). These platforms provide open-source data that is accessible to anyone through their respective websites.
Table 2: REVISE: Data Attributes and Preprocessing for Case Study of Seoul
N Urban Attribute Abbreviation Data Source Acquisition Process
1 Housing Density HOUS_DEN Seoul Housing Statistics by Neighborhood Areal weighting interpolation using census boundaries
2 Bus Stop Density BUS_ST_DEN Smart Card Transit Location Database Spatial density calculation with 500m grid aggregation
3 Floating Population FL_POP_DEN Seoul Hourly Population Statistics Mobile signaling data aggregation (8AM-8PM average)
4 Foreign Population FOR_POP_DEN Seoul Short-term Resident Population Data Immigration record filtering and temporal alignment
5 Commuting Population CMM_POP_RATIO Seoul Mobility Pattern Survey Morning commute ratio calculation (7-9AM)
6 Incoming Population I_POP_RATIO Seoul Mobility Pattern Survey Peak hour inflow-outflow ratio computation
7 Parking Space Density PS_DEN Kakao Map POI Database POI classification and kernel density estimation
8 Metro Coverage METRO_COV_P Seoul Metro Station Master Data 1,000m buffer analysis around station centroids
9 Disabled Population DIS_POP_P Seoul Disability Registry Population proportion calculation per administrative district
10 Beneficiary BEN_POP_P Seoul Basic Livelihood Security Registry Welfare recipient percentage computation
11 Senior Population SEN_POP_P Seoul Elderly Population Statistics Age-based demographic proportion analysis
12 Average Income AVG_INC Seoul Commercial District Analysis System Neighborhood-level income aggregation
13 Touristic Facility Density TOUR_DEN Kakao Map POI Database Tourism facility identification and density mapping
14 Cultural Facility Density CULT_DEN Kakao Map POI Database Cultural facility identification and density mapping
15 Energy Usage Intensity EN_USE_INT Seoul Energy Consumption Report Annual kWh normalization by land area
16 Employment Density EMPL_DEN Seoul Business Establishment Registry Workforce density calculation per km²
17 Bus Users Per Stop BUS_USR_DEN Seoul Public Transit Ridership Data Boarding counts normalized by stop locations
18 Metro Users Per Station METRO_USR_DEN Seoul Metro Ridership Statistics Station-level usage intensity calculation
19 Spending Ratio SPND_P Seoul Commercial District Analysis System Expenditure-to-income ratio computation
20 Land Price LND_PR Seoul Official Land Price Index Parcel value aggregation per administrative district
21 Transportation Expenditure TR_EXP_P Seoul Commercial District Analysis System Transit cost percentage calculation from income data
22 Distance to Helipads HELI_DIST National Aviation Facility Registry Euclidean distance from census tract centroids
23 Distance to Hospitals HOSP_DIST National Medical Institution Database Straight-line distance calculation
24 Distance to VFR Routes VFR_DIST Aeronautical Navigation Chart System Air corridor buffer analysis
25 Distance to Fire Stations FS_DIST Emergency Service Location Database Service area proximity modeling
Note: All metrics at census tract level. VFR = Visual Flight Rules. GIS processing applied.
The descriptive statistics table (Table 3) summarize 426 observations across 25 urban attributes selected for vertiport site selection in Seoul. The data reveals substantial variability in urban characteristics, with several metrics showing notable statistical properties. Population-related measures (e.g., FL_POP_DEN, FOR_POP_DEN) and specialized facilities (e.g., CULT_DEN) exhibit right-skewed distributions, indicating concentrated density in specific areas. Infrastructure coverage metrics like METRO_COV_P demonstrate more uniform distribution across the city. These distribution patterns highlight the heterogeneity of Seoul’s urban landscape, with implications for optimal vertiport placement. The data’s non-normal distribution characteristics informed our subsequent analytical approach using principal axis factoring rather than standard methods requiring normality assumptions.
Table 3: Descriptive Statistics of Raw Urban Attributes
Variable Count Mean Std Min Median Max Skewness Kurtosis
HOUS_DEN 426 7109.123 3811.969 0.000 6824.086 20192.857 0.407 -0.001
FL_POP_DEN 426 23216.819 12506.621 0.000 22167.694 100436.473 1.011 3.210
FOR_POP_DEN 426 347.584 1061.019 0.000 83.023 13992.351 8.002 81.075
CMM_POP_RATIO 426 4445.137 2485.035 0.000 4058.770 15616.181 0.784 0.960
I_POP_RATIO 426 0.970 0.833 0.000 0.744 7.033 3.741 18.591
DIS_POP_P 426 0.041 0.017 0.000 0.040 0.174 2.498 14.291
BEN_POP_P 426 0.040 0.031 0.000 0.036 0.236 2.286 9.521
SEN_POP_P 426 0.182 0.040 0.000 0.181 0.359 0.195 1.837
AVG_INC 426 3.308 1.099 0.000 3.101 7.421 0.249 2.164
SPEND_P 426 2128.242 10895.023 0.000 477.260 140851.873 9.826 105.405
TR_EXP_P 426 0.067 0.107 0.000 0.026 0.908 3.449 16.445
BUS_ST_DEN 426 24.289 12.967 0.000 22.518 84.783 1.003 1.639
METRO_COV_P 426 0.662 0.319 0.000 0.754 1.000 -0.651 -0.862
METRO_USR_DEN 426 5538.780 11683.466 0.000 7351.026 110019.000 3.640 20.021
BUS_USR_DEN 426 465.426 268.217 0.000 386.614 1397.289 1.064 0.686
PS_DEN 426 9794.966 5851.015 0.000 9363.484 70584.091 3.041 26.676
TOUR_DEN 426 1.620 2.489 0.000 1.102 28.085 5.045 39.433
CULT_DEN 426 4.636 13.141 0.000 1.588 206.410 9.905 135.248
EN_USE_INT 426 1.000 1.001 0.032 0.688 8.540 3.826 20.183
EMPL_DEN 426 11269.860 12975.330 0.000 7248.035 109586.870 3.378 15.490
LAND_PR 426 4.663 3.092 0.000 3.802 23.770 2.284 6.917
HELI_DIST 426 1.946 1.486 0.013 1.534 8.109 1.187 1.274
HOSP_DIST 426 3.421 2.073 0.040 2.931 10.524 0.987 0.727
VFR_DIST 426 3.103 2.169 0.005 2.808 8.751 0.537 -0.667
FS_DIST 426 2.037 1.143 0.000 1.846 7.244 0.959 1.949
4.2 LFA Results
Based on Kaiser’s criterion, six factors were retained from the 25 urban attributes, collectively accounting for 66.82% of the total cumulative variance.
Table 4: Resultant Factors, Eigenvalues, and Variances from LFA
Factor Initial Eigenvalues % of Variance Cumulative Variance
1 3.405219 13.620877 13.620877
2 3.847866 15.391465 29.012342
3 2.946635 11.786539 40.798881
4 2.506135 10.024542 50.823423
5 2.742083 10.968330 61.791753
6 1.256834 5.027335 66.819088
Table 4 details the eigenvalue of each factor, representing the variance in the original variables explained by that factor. The first factor had an eigenvalue of 3.41, accounting for 13.62% of the total variance. Factor 2 had the highest eigenvalue of 3.85, which explained 15.39% of the total variance. In total, the six identified factors explain 66.82% of the total variance. These results indicate that a relatively small set of underlying factors can account for a substantial portion of the variance in the original attributes, supporting their use for further analysis and interpretation.
For each attribute, the factor with the highest absolute factor loading was identified, designating the attribute as a “dominant attribute” for that factor. These dominant attributes, grouped within the same factor, highlight shared characteristics that define the factor (see Figure 2). Based on the overarching characteristics, each factor was interpreted (see Table 5). The impact directions were inferred on the basis of the associated variables and their anticipated influences on the vertiport placement.
Figure 2: Structural LFA results showing the dominant attributes for each latent factor.
Table 5: Factor interpretations and factor loadings
Latent Factors Impact Dominant Attributes Factor Loadings Eigenvalue % Of Variance
1. Economic Dynamism (+) EN_USE_INT 0.747 3.41 13.62
EMPL_DEN 0.729
METRO_USR_DEN 0.656
SPND_P 0.638
LND_PR 0.522
2. Community Preparedness (-) HOUS_DEN 0.785 3.85 15.39
BUS_ST_DEN 0.708
FL_POP_DEN 0.807
PS_DEN 0.756
METRO_COV_P 0.793
3. Societal Equity (+) DIS_POP_P 0.908 2.95 11.79
BEN_POP_P 0.896
SEN_POP_P 0.800
AVG_INC -0.528
4. Operation Preparedness (-) HELI_DIST 0.758 2.51 10.02
HOSP_DIST 0.714
VFR_DIST 0.524
FS_DIST 0.435
5. Urban Vibrancy (+) TOUR_DEN 0.758 2.74 10.97
CULT_DEN 0.759
I_POP_RATIO 0.654
FOR_POP_DEN 0.630
6. Mobility Patterns (+) BUS_USR_DEN 0.724 1.26 5.03
TR_EXP_P -0.511
4.3 Identified Factors into Suitability and Attractiveness
Based on the LFA results, the framework establishes two essential composite measures: suitability (derived from factors 2, 3, and 4) and attractiveness (from factors 1, 5, and 6). These complementary metrics capture different aspects of vertiport viability - suitability addresses operational feasibility considerations, while attractiveness reflects operational merit indicators.
Suitability Factors
Factor 2: Community Preparedness encompasses the density of housing units, bus stations, floating population, car parking spaces, and metro service coverage. This factor receives a negative weight to account for potential community resistance and challenges in the initial adoption of UAM.
Concerns about ground population safety near vertiports have been noted by Yedavalli and Mooberry (2019), highlighting potential resistance from communities. Oh and Yoon (2022) stressed the importance of accounting for spatio-temporal variations in population density to assess and mitigate risks associated with UAS operations in urban areas. Additionally, Smith et al. (2024) emphasized the importance of public engagement and transparent communication about UAM’s impacts to foster acceptance. Similar to the adverse effects of air transportation noise (Friedt and Cohen, 2021), frequent UAM operations could generate significant noise pollution, making densely populated areas less suitable for vertiport placement. Oh and Yoon (2024) highlighted the need to balance operational feasibility with minimizing risks in densely populated regions through a Pareto-based framework for urban airspace assessment.
While established transit infrastructure may offer connectivity advantages (Straubinger et al., 2020; Wang et al., 2023; Yan et al., 2024), competition with existing transportation modes (Hasan, 2019), induced traffic (Yadav et al., 2025), and safety risks (Al Haddad et al., 2020) could pose significant concerns during early UAM adoption--particularly in cities like Seoul. These considerations suggest that areas with high density and extensive transit connectivity may be less suitable for vertiport placement.
Factor 3: Societal Equity represents the demographic and socioeconomic attributes of metropolitan areas, including the ratios of disabled populations, national basic livelihood recipients, senior populations, and average income levels. These indicators help identify regions with concentrations of vulnerable populations, offering essential insights for suitability assessments that prioritize social justice.
Factor 4: Operation Preparedness is formulated based on proximity to critical emergency infrastructures such as fire stations, major hospitals, helipads, and VFR routes. These variables reflect a region’s preparedness to meet operational requirements and support emergency responses. In contingency scenarios at vertiports, proximity to these critical facilities ensures faster reaction times, enhancing safety—particularly important for mitigating risks associated with battery-operated UAM aircraft (Wisk, 2022).
Additionally, close access to emergency services improves response capabilities in traffic-congested areas, where UAM vehicles can serve as vital reinforcements to the emergency response network. Existing helipads and VFR routes provide foundational infrastructure that can be leveraged during the initial stage of UAM implementation, further supporting operational suitability (Federal Aviation Administration, 2020).
Attractiveness Factors
Factor 1: Economic Dynamism is gauged through energy consumption relative to the average, working population density, weekday hourly metro users per station, spending-to-income ratio, and official land prices. These metrics collectively reflect the economic vibrancy of an area, establishing a foundation for evaluating attractiveness. This aspect is particularly important given the relatively high expected cost of UAM flights (\$3.50 to \$4.00 per passenger-km) during the initial stages of implementation (Binder et al., 2018; Rimjha et al., 2021).
Factor 5: Urban Vibrancy evaluates the centrality of regions within Seoul, measured by the density of touristic and cultural facilities, the presence of foreigners, and the ratios of moving and incoming populations. These variables together indicate the attractiveness of regions by revealing their centrality and dynamism.
Factor 6: Mobility Patterns are assessed through weekday hourly bus users per stop and the inversely related ratio of transportation expenditures to total spending. Areas with high bus usage combined with low transportation spending suggest either higher reliance on personal vehicles or higher incomes that make public transportation expenses relatively insignificant. These regions with established travel demand and lower relative transportation costs indicate potential attractiveness for adopting new transportation modes like UAM.
4.4 Suitability and Attractiveness Scores
The framework categorizes urban factors into two dimensions critical for UAM site selection: suitability (operational feasibility) and attractiveness (operational merits). Figure 3 illustrates their spatial distribution across Seoul, revealing distinct geographic patterns. Central business districts exhibit concentrated attractiveness scores, reflecting their inherent advantages in demand generation and economic activity. In contrast, suitability scores display broader dispersion, with elevated values observed in less congested peripheral neighborhoods. This dichotomy establishes a foundational tension between immediate operational viability and long-term scalability in vertiport planning.
Figure 3: Spatial Distribution of Suitability and Attractiveness Scores in Seoul
5 Result Analysis and Discussions
5.1 Comparative Analysis of Top-Scoring Regions
Figure 4 highlights the contrasting characteristics of Seoul’s top 10 regions for suitability and attractiveness. Suitability leaders demonstrate uniform score distributions from Dobong 1-dong (Suitability Score: 5.19) to Ganggil-dong (4.06) making all regions stable candidates for phased UAM deployment. Conversely, attractiveness scores show extreme spatial skewness, dominated by central hubs like Myeong-dong (Attractiveness Score: 12.81) and Jongno 1.2.3.4-dong (8.81). This divergence implies the distinct implementation strategies: concentrated initial deployments in high-attractiveness zones versus distributed networks prioritizing suitability.
Figure 4: Top 10 Regions for Suitability (blue-colored) and Attractiveness Scores (red-colored)
5.2 Factor Decomposition of Top-Scoring Regions
Figure 5 reveals the latent factor compositions driving regional scores. When percentage contributions are examined, the high-suitability regions (left cluster) derive 46.7% of their scores from Factor 2 (Community Preparedness) and Factor 3 (Societal Equity) on average, emphasizing dominant contribution of these latent factors on suitability. In contrast, top-attractiveness regions (right cluster) rely predominantly (63.3%) on Factor 5 (Urban Vibrancy) and Factor 1 (Economic Dynamism), with Yeoksam-1-dong showing exceptional Factor 1 dominance (64.6%).
Figure 5: Factor scores for the top 10 suitability and attractiveness regions with absolute contribution in percentages
The figure highlights that suitability leaders exhibit significant positive contributions from Factor 2 (mean 20.3%), emphasizing its critical role in determining regional suitability. In contrast, Factor 2 scores are mostly negative or negligible in regions with high attractiveness. Conversely, Factor 5 demonstrates the opposite trend, further underscoring key differences between suitability and attractiveness. These findings suggest that community preparedness and urban vibrancy within the urban environment act as pivotal discriminators between suitability and attractiveness.
Insights derived from Seoul’s analysis can be extended to scenarios where regional data is scarce or difficult to obtain. Without the need to implement the full USE-LFA framework, decision-makers in cities with urban environments similar to Seoul can leverage community preparedness and urban vibrancy as proxies to approximate regional suitability and attractiveness. This demonstrates the potential applicability of the USE-LFA framework in cities with limited data availability.
5.3 Quadrant Map of Suitability and Attractiveness
The quadrant map in Figure 6 visualizes the spatial distribution of regions in Seoul based on their suitability and attractiveness scores. By defining median thresholds for both metrics, the map categorizes regions into four quadrants, revealing distinct spatial typologies and clustering patterns.
Figure 6: Quadrant Map of Suitability and Attractiveness Scores in Seoul (Magenta: Attractiveness-biased; Cyan: Suitability-biased; Green: Balanced Regions)
Three primary typologies emerge. Balanced regions shown in green triangles represent areas with high scores for both suitability and attractiveness, with minimal differences between the two scores. Examples include Namhyeon-dong, Namyeong-dong, and Songcheon-dong. These regions are likely to offer comprehensive benefits due to their equilibrium between urban qualities. Suitability-biased regions in cyan squares exhibit the largest score differences, with suitability scores dominating. Examples include Dobong 1-dong, Beon 3-dong, and Gayang 2-dong. Conversely, attractiveness-biased regions drawn in magenta circles, such as Myeong-dong, Sogong-dong, and Yeoksam 1-dong, demonstrate exceptionally high attractiveness scores with significant differences compared to their suitability scores.
The broader distribution of data points reveals several notable patterns. (1) Most regions cluster near median values, indicating moderate performance in both metrics. (2) An elliptical cloud stretching from lower left to upper right suggests a loose positive correlation between suitability and attractiveness. (3) The lower left quadrant highlights under-performing regions that may require targeted urban renewal efforts. (4) High-density regions near the origin (median intersection) represent typical Seoul regions with average urban qualities. These typologies highlight Seoul’s urban configuration while offering valuable insights for policy applications. Cities with similar quadrant distributions could adopt transferable strategies tailored to their specific typologies.
An intriguing observation is that regions with extreme biases in either suitability or attractiveness are not positioned near the lowest values of the opposite metric. Instead, they tend to cluster around the median point of the opposite scoring group. This suggests that regions with pronounced strengths in one dimension still maintain moderate performance in the other dimension rather than being entirely deficient.
While the quadrant map provides intuitive visual insights into regional typologies, its reliance on subjective thresholds may limit its applicability in scenarios requiring adaptable prioritization between suitability and attractiveness. To address this limitation and enhance decision-making flexibility, we introduce the concept of the v-score.
5.4 v-score
The v-score is a composite metric that combines suitability and attractiveness scores using a weighting parameter $\alpha$. This parameter, which ranges from 0 to 1, allows users to adjust the relative importance of suitability and attractiveness based on preference states or strategy-specific priorities. The formulation of the v-score confines suitability and attractiveness scores as follows:
[…]
5.4.1 Spatial Sensitivity to Different Weighting Parameter
When implementing different prioritization parameters, clear spatial variations emerge. Figure 7 illustrates how the distribution of high-scoring regions shifts as $\alpha$ increases from 0 (full attractiveness priority) to 1 (full suitability priority).
Figure 7: Spatial distribution of the top 30 regions with the highest v-scores under varying levels of suitability prioritization ($\alpha$).
At $\alpha=0$, where attractiveness is fully prioritized, high v-score regions cluster in affluent, economically prominent areas of Seoul, reflecting the strong link between attractiveness and established economic centers. As $\alpha$ increases to 0.5, representing an equal prioritization of suitability and attractiveness, the distribution shifts, with high v-score regions becoming more dispersed and extending towards the outer parts of Seoul. At $\alpha=1$, fully prioritizing suitability, high v-score clusters reappear, primarily in the norther and western parts of Seoul. These transitions demonstrate how prioritization significantly influences optimal site selection, suggesting that balanced approaches may identify different locations than those focusing solely on market potential or operational requirements. To further analyze the impact of different prioritization, the convex combinations were examined with the contour plots of the v-scores.
5.4.2 Convex Combination of Suitability and Attractiveness
Figure 8: Selected regions in red dots with the v-score higher than the thresholds 2 (from a to c) and 3 (from d to f)
The analysis of the convex combination of suitability and attractiveness scores reveals significant insights into potential vertiport site selection. Three different weights were examined, with the slope of the v-score contours varying as the weight for suitability increased. This demonstrates the growing influence of suitability on the composite v-score as its weight increases. When applying a v-score threshold of $\theta>2$, the number of identified regions changes notably. At weight 0 (37 selected regions), significantly fewer regions meet the criteria than at weight 1 (55 regions). This disparity highlights the skewed nature of attractiveness scores and suggests that fewer regions excel in attractiveness compared to suitability.
5.4.3 Threshold-Dependent Weighting Effects
Analysis of scoring thresholds and weights revealed distinct patterns in the number of candidate regions. Figure 9 shows the relationship between preference weight ($\alpha$) and region counts across seven threshold values ($\theta$). All curves exhibit a convex shape: region counts initially decrease as $\alpha$ increases from 0, reach a minimum at intermediate $\alpha$ values (typically 0.2-0.6), then increase as $\alpha$ approaches 1.0. This pattern indicates that balanced weighting creates the most stringent filtering effect, suggesting regions excelling in both dimensions are less common than those performing well in just one.
Figure 9: illustrates the relationship between preference weight $\alpha$ and the count of regions across seven different $\theta$ values.
Table 6 quantifies these relationships, showing region counts and percentages out of 426 total regions in Seoul for different threshold-weight combinations. Sensitivity to weighting is more pronounced at lower thresholds ($\theta\leq 1.5$), where the changes curves are steeper and the percentage difference between the maximum and minimum number is greater. At higher thresholds, this sensitivity diminishes. The minimum points on most curves occur more during the lower values of the weight (around $\alpha$ = 0.4), indicating suitability is somewhat more stringent measure than attractiveness.
Threshold ($\theta$) Weight ($\alpha$)
0.0 0.2 0.4 0.6 0.8 1.0
1.0 76 (17.8%) 72 (16.9%) 68 (16.0%) 83 (19.5%) 95 (22.3%) 106 (24.9%)
1.5 55 (12.9%) 36 (8.5%) 40 (9.4%) 54 (12.7%) 64 (15.0%) 78 (18.3%)
2.0 37 (8.7%) 27 (6.3%) 26 (6.1%) 31 (7.3%) 44 (10.3%) 55 (12.9%)
2.5 27 (6.3%) 20 (4.7%) 14 (3.3%) 15 (3.5%) 25 (5.9%) 38 (8.9%)
3.0 20 (4.7%) 15 (3.5%) 8 (1.9%) 7 (1.6%) 13 (3.1%) 27 (6.3%)
3.5 16 (3.8%) 12 (2.8%) 6 (1.4%) 2 (0.5%) 8 (1.9%) 16 (3.8%)
4.0 14 (3.3%) 7 (1.6%) 4 (0.9%) 2 (0.5%) 3 (0.7%) 10 (2.3%)
Table 6: Counts and percentages (out of 426 regions), step increment of $\alpha=0.2$.
These findings demonstrate the critical role of strategic weighting decisions in reshaping candidate region selection. The observed convex patterns and threshold sensitivity highlight the necessity for stakeholders to explicitly prioritize either suitability or attractiveness during site selection processes. Balanced weighting ($\alpha=0.4-0.6$) emerges as a stringent filter, favoring regions that excel in both dimensions, while extreme weight values emphasize specialization in one dimension over the other. This insight provides a foundation for informed decision-making, ensuring that site selection strategies align with strategic objectives.
6 Conclusion
This study presents the Urban Site Evaluation with Latent Factor Analysis (USE-LFA) framework, a comprehensive, data-driven methodology for evaluating urban port sites, demonstrated through a case study of vertiport site selection in Seoul. By integrating Latent Factor Analysis (LFA), the framework identifies six key urban factors-Economic Dynamism, Community Preparedness, Societal Equity, Operational Preparedness, Urban Vibrancy, and Mobility Patterns-categorized into two dimensions: Suitability and Attractiveness. These enable a dual-perspective evaluation of urban environments, balancing operational feasibility and operational merit. The framework’s adaptability is underscored by the development of a composite v-score, which allows stakeholders to prioritize suitability or attractiveness based on strategic objectives. The analysis revealed how optimal site selection shifts spatially depending on prioritization, highlighting the tension between centralized economic hubs and more dispersed operationally viable regions. The quadrant-based classification of regions and convex combination analysis further provide informative insights for tailoring urban port integration strategies.
Unlike existing tools such as NASA’s VAMOS! (Sheth, 2023), which often conflate operational feasibility and merits into a single evaluation, USE-LFA distinctly separates these dimensions while integrating broader urban environmental factors such as community preparedness and societal equity. This separation enhances the framework’s versatility, allowing it to adapt to diverse urban contexts and enabling stakeholders to evaluate suitability and attractiveness independently while tailoring decisions to specific regional needs. Furthermore, its flexible design supports applications beyond vertiports, such as transit hubs and other emerging mobility infrastructures. By systematically addressing the interplay between operational feasibility and metrits, USE-LFA provides informative insights for the systematic integration of ports into complex urban environments. Policymakers can leverage this framework to align port placement with strategic goals such as improving accessibility in underserved areas or enhancing emergency response capabilities. Urban planners can use factor-specific insights to design infrastructure that complements existing systems while addressing community concerns. Moreover, the framework’s flexibility allows for continuous adaptation to evolving urban environments or new data inputs.
Future research can expand our framework by incorporating additional layers such as noise impact assessments (Tan et al., 2023, 2024), airspace constraints, and dynamic urban changes. For example, integrating optimized noise-aware flight trajectories with regional characteristics could support sustainable UAM adoption (Ho-Huu et al., 2019; Gao et al., 2023). By further addressing these components, our framework could be utilized to ensure that UAM systems remain equitable and efficient across diverse contexts.
Acknowledgements
This work is financially supported by Korea Ministry of Land, Infrastructure, and Transport (MOLIT) as Innovative Talent Education Program for Smart City and by the National Research Foundation of Korea(NRF) grant funded by the Korea government (MSIT) (No. 2020R1A2C2010200).
Generative AI and AI-assisted Technologies
During the preparation of this manuscript, the authors used ChatGPT and Grammarly to check the grammar. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Authors Contributions
Sungmin Sohn: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing – Original Draft, Visualization, Project administration. Namwoo Kim: Conceptualization, Methodology, Validation, Resources, Supervision. Mark Hansen: Validation, Investigation, Supervision. Yoonjin Yoon: Conceptualization, Methodology, Validation, Resources, Writing – Review & Editing, Supervision, Project administration, Funding acquisition.
Appendix A Notation Table with Descriptions
The following notation establishes the fundamental components of our urban latent factor analysis for vertiport site selection (Table 7).
Table 7: Definition of Notation for Latent Factor Analysis
Symbol Dimensions Description
[…]
$\mathbf{A}=[a_{ij}]$ $\mathbb{R}^{N\times R}$ Standardized urban attribute matrix; $a_{ij}$ is the standardized measure of attribute $i$ in region $j$.
$\mathbf{L}=[\ell_{im}]$ $\mathbb{R}^{N\times M}$ Urban attribute loading matrix; $\ell_{im}$ represents the intensity of relationship between attribute $i$ and latent factor $m$.
$\mathbf{F}=[f_{mj}]$ $\mathbb{R}^{R\times M}$ Urban factor score matrix; $f_{mj}$ indicates the degree to which region $j$ embodies latent factor $m$.
$\mathbf{E}$ $\mathbb{R}^{N\times R}$ Error matrix capturing region-specific variance unique to each attribute.
$\mathbf{R}$ $\mathbb{R}^{N\times N}$ Correlation matrix of raw urban attributes.
$\hat{\mathbf{c}}_{i}^{2(0)}$ $\mathbb{R}^{+}$ Initial communality for attribute $i$, calculated as $1-\frac{1}{r^{ii}}$ where $r^{ii}$ is the $i^{\text{th}}$ diagonal element of $\mathbf{R}^{-1}$.
$\mathbf{R}^{*}$ $\mathbb{R}^{N\times N}$ Adjusted correlation matrix with diagonal elements replaced by communality estimates.
$\mathbf{Q}$ $\mathbb{R}^{N\times N}$ Eigenvector matrix from decomposition of $\mathbf{R}^{*}$; columns are eigenvectors $\boldsymbol{q}_{m}$.
$\mathbf{\Lambda}$ $\mathbb{R}^{N\times N}$ Diagonal matrix of eigenvalues $\lambda_{1},...,\lambda_{N}$ from decomposition of $\mathbf{R}^{*}$.
$\mathbf{L}^{(k)}$ $\mathbb{R}^{N\times M}$ Factor loading matrix at iteration $k$, computed from eigenvectors and eigenvalues.
$\hat{\mathbf{c}}_{i}^{2(k+1)}$ $\mathbb{R}^{+}$ Updated communality for attribute $i$ at iteration $k+1$, calculated as $\sum_{m=1}^{M}(\ell_{im}^{(k)})^{2}$.
$\epsilon$ $\mathbb{R}^{+}$ Convergence threshold ($10^{-5}$) for terminating iterative updates.
$\mathbf{L}^{*}$ $\mathbb{R}^{N\times M}$ Final factor loading matrix containing $M$ factors that satisfy Kaiser criterion ($\lambda_{m}\geq 1$).
$\mathbf{L}^{R}$ $\mathbb{R}^{N\times M}$ Rotated factor loading matrix after Varimax rotation for improved interpretability.
$\mathbf{T}$ $\mathbb{R}^{M\times M}$ Orthogonal rotation matrix satisfying $\mathbf{T}^{\top}\mathbf{T}=\mathbf{I}_{M}$.
$\mathbf{B}=[\beta_{mi}]$ $\mathbb{R}^{M\times N}$ Factor scoring weights calculated as $(\mathbf{L}^{\top}\mathbf{R}^{-1}\mathbf{L})^{-1}\mathbf{L}^{\top}\mathbf{R}^{-1}$.
$\mathbf{T}$$\{0,1\}^{2\times M}$ Composite definition matrix; $t_{1m}=1$ if factor $m$ belongs to Suitability, $t_{2m}=1$ if Attractiveness.
$\hat{\mathbf{F}}$ $\mathbb{R}^{2\times R}$ Composite score matrix computed as $\mathbf{T}\mathbf{F}$, containing Suitability and Attractiveness scores for each region.
$\mathcal{S}$$\subset\{1,...,M\}$ Set of indices for Suitability factors.
$\mathcal{A}$$\subset\{1,...,M\}$ Set of indices for Attractiveness factors, where $\mathcal{S}\cap\mathcal{A}=\emptyset$.